How many 3 inch tiles to fill an area that is 216 sq ft?

Answer by checkley79(3341)

IF THE TILE IS A 3 INCH SQUARE THEN:
3^3=9 IN^2 OR
12*12/9=144/9=16 SQUARES PER SQUARE FOOT.
216^2/16
46,656/16=2,916 TILES ARE NEEDED.